OpenStudy (anonymous):

Given the triangle and equation below, what is the value of x? sec∠C - cot∠A = 3x

OpenStudy (anonymous):



OpenStudy (amistre64):

sec∠C - cot∠A = 3x cotA = tanC secC - tanC = 3x 1 - sinC = 3x cosC 1 - sinC = 3x (6/7) seems like either way you go youll need to determine ab

OpenStudy (jdoe0001):

\(c = \sqrt{7^2+6^2}\)

OpenStudy (jdoe0001):

wait, that.... didn't go well :/

OpenStudy (mathstudent55):

1. Use the Pythagorean theorem to find AB 2. secC = 1/cosC 3. cotA = 1/tanA 4. Use SOH CAH TOA and the identities above to find secC and cotA. 5. Set up the equation and solve for x.

OpenStudy (amistre64):

lol, confusing names with convention are ya

OpenStudy (jdoe0001):

we need the opposite, thus \(b = \sqrt{7^2-6^2}\)

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

ok thanks guys but can u choose the multiple choice

OpenStudy (anonymous):

idk how to use the SOH CAH TOA to find secC and tanA

OpenStudy (mathstudent55):

SOH: sin = opp/hyp CAH: cos = adj/hyp TOA: tan = opp/adj There is no cot or sec listed, but cot = 1/tan, and sec = 1/cos, so cot = adj/opp and sec = hyp/adj.

OpenStudy (anonymous):

ok i understand now thanks so i narrowed it down to either being A or C im just gonna take it from here thanks alot

OpenStudy (mathstudent55):

wlcm